The pH of a solution by mixing 100 ml of 0.2M CH3COOH with 100 ml of 0.2M NaOH. pKa for CH3COOH = 4.74

Grade:11

1 Answers

Vikas TU
7 years ago
               CH3COOH + H2O ----------> CH3COO-   +       H3O+
At eqm.      0.2-x                                  x                       x
for x very very less 0.2 – x = 0.2
Hence, x^2/0.2 = anitlog(Ka)
x = 0.001908 approx.
[H3O+] = 0.00198
and 
OH- = > (0.02-0.00198)/2 = > 0.099 approx. 
Hence, pH = -log(1.01x10^-13) => 12.99 or 13.